a scientist develops a formula which can turn a light bulb into a camel.working alone it takes the scientist 4 hours to create one dose of formula.the scientist assistant can make one dose in 12 hours. how long working together will it take the scientist and his assistant to create one dose of formula? please answer the question by using a table and the 5 step method. thank you.
Answer by stanbon(75887)

a scientist develops a formula which can turn a light bulb into a camel.
---------
working alone it takes the scientist 4 hours to create one dose of formula.
---------
the scientist assistant can make one dose in 12 hours.
---------
how long working together will it take the scientist and his assistant to create one dose of formula?
---
Your able has three rows: scientist ; assistant ; together
---
Your table has two coluns: time and rate
-----
Scientist time = 4 hrs ; rate = 1/4 job/hr
---
Assistant time = 12 hrs ; rate = 1/12 job/hr
---
Together time = x hrs ; rate = 1/x job/hr
-------------------------------------------
Equation:
rate + rate = rate
1/4 + 1/12 = 1/x
Multiply thru by 12x
---
3x + x = 12
4x = 12
x = 3 hrs (time to do the job together)
=========
Cheers,
Stan H.